OpenStudy (anonymous):

Will give medal to most helpful!!! What is the slope of the line 4y-6x+12=0 in simplest form?

OpenStudy (anonymous):

Just switch it to slope-intercept form and look at slope? Can you do that?

OpenStudy (anonymous):

hmmmm, is it y=6x+4?

OpenStudy (anonymous):

so the slope would be 6?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 4y - 6x + 12 = 0 move x and 12 to other side - 4y = 6x - 12 divide by 4 to isolate y - y = 6/4x -3 so slope is 6/4 in simplest form = 3/2

OpenStudy (anonymous):

ohhhhhhhhhhhh ok!thank you SO much:D

OpenStudy (anonymous):

of course happy to help